OpenStudy (ray1998):

Please help! Will give medal and fan. Thank you in advance :) 1. Simplify the expression. -5 + i/2i

OpenStudy (ybarrap):

1/i = -i So i/2i = i/2 * -i = \(\cfrac{-i^2}{2}\) See that? Then -5 + i/2i = \(-5+\cfrac{-i^2}{2}\) Then \(i^2=-1\) Can you finish it?

OpenStudy (ray1998):

Okay. Thank you. I don't think I can honestly, I'm not at all familiar with these problems. I'm very behind in my Algebra class. I don't expect you to give me the answer, but if you could just work with me on it :) that would be great. @ybarrap

OpenStudy (ybarrap):

You have $$ -5 +\cfrac{i}{2i} $$ Right?

OpenStudy (ray1998):

Right

OpenStudy (ybarrap):

I explained that 1/i = -i, which you can use; however you can also cancel the i's: $$ -5 +\cfrac{\cancel i}{2\cancel i}=-5+\cfrac{1}{2} $$ See that?

OpenStudy (ybarrap):

This is the easier approach; however, the 1st approach shows how things work if you play with the imaginary number, i. BTW, \(i=\sqrt{-1}\) so \(i^2=-1\), just in case

OpenStudy (ybarrap):

does any of this make any sense?

OpenStudy (ray1998):

Okay. So it's easiest to keep the i's in the equation?

OpenStudy (ray1998):

Em, I'm trying to grasp it all lol.

OpenStudy (ybarrap):

yes, if you cancel them, you can just deal with them like regular old numbers

OpenStudy (ybarrap):

keep trying, it's fulfilling and exciting

OpenStudy (ray1998):

Okay. Lol I'm not so sure about it being exciting, so much as complicated.

OpenStudy (ybarrap):

did I say that, lol

OpenStudy (ray1998):

Yes lol

OpenStudy (ybarrap):

There was a time when people thought that you could not take a square root of a negative number, then someone decided, not that complicated, just make it the letter "i" and deal with it like an ordinary number and see what happens.

OpenStudy (ray1998):

Okay.

OpenStudy (ybarrap):

So once you've simplified you have $$ -5 + \cfrac{1}{2}=-4\cfrac{1}{2}=-\cfrac{9}{2} $$

OpenStudy (ray1998):

Alright

OpenStudy (ybarrap):

good luck \(:-) \)

OpenStudy (ray1998):

Thank you @ybarrap

OpenStudy (ybarrap):

You're welcome